qwevery:

A pair of shoes usually sells for $62. If the shoes are 40% off, and sales tax is 5%, what is the total price of the shoes, including tax? A. $42.32 B. $39.06 C. $40.30 D. $35.81

kbear:

62 - .40 × 62 = 37.20 37.20 + 37.20 × .08 = 40.18

qwevery:

thats not up there-

SmokeyBrown:

kbear has the right idea, but I think they typed in the sales tax wrong. If the tax were 8% instead of 5%, they'd be correct. Picking up from where they left off... After the 40% discount, the shoes would be 37.20. Now let's apply the 5% tax: 37.20 * 1.05 = 39.06, which should be the final price.